OpenStudy (anonymous):

What is f(-3) for the function f(x) = -5x - 7

jimthompson5910 (jim_thompson5910):

Replace every copy of x with -3, then evaluate f(x) = -5x - 7 f(-3) = -5(-3) - 7 f(-3) = 15 - 7 f(-3) = 8